can lattice energy be estimated by born-haber cycle? what about hydration energy and electron gain enthalpy?

Grade:12

1 Answers

SAGAR SINGH - IIT DELHI
13 years ago

Dear student,

The answer is yes.

Lattice energy of ionic solids

Born Haber Cycle helps us to calculate the lattice energy of ionic solid, provided other thermodynamic data is known. For example, the lattice energy of magnesium fluoride (MgF2) can be calculated when the sublimation energy (S) of Mg = 146.4 kJ mol-1; IE1 and IE2 values of Mg=737 and 1449 kJ mol -1 respectively; Dissociation energy (D) of fluorine = 158.8 kJ mol -1; EA of fluorine = - 328 kJ mol -1 and DHf of MgF2 =- 1096.5 kJ mol 1. Born Haber Cycle for MgF2 is as shown:

Born Haber Cycle for MgF2

Born Haber Cycle for MgF2

DHf = DHs + DHd + (IE1 + IE2) + 2 EA1+ U

U = DHf - DHs - DHd - (IE1 + IE2) - 2 EA1

= -1096.5 - 146.4 - 158.8 - (737 + 1449) - 2(-328)

= -2931.7 kJmol-1

Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

 

Askiitians Expert

Sagar Singh

B.Tech, IIT Delhi

sagarsingh24.iitd@gmail.com